OpenStudy (anonymous):

SOLVE THE LOGARITHMIC EQUATION LOG3(X+6)+LOG3(X-6)-LOG3X=2

OpenStudy (anonymous):

I know most but am stuck

OpenStudy (anonymous):

Do you know log properties, or do you know exponent properties?

OpenStudy (anonymous):

yes . log3 (x+6)(x-6)/x=2

OpenStudy (anonymous):

Okay, so do you know the inverse of a log function?

OpenStudy (anonymous):

would it be 3^2=(x+6)(x-6)/x

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

well thanks you helped me see my error x=12

OpenStudy (anonymous):

There are two solutions

OpenStudy (anonymous):

Wait, nevermind

OpenStudy (anonymous):

I guess technically the negative solution doesn't count because the original equation makes it so \(x>0\).